# S1-Bench: A Simple Benchmark for Evaluating System 1 Thinking Capability of Large Reasoning Models

### v1

@@ Abstract
We introduce S1-Bench, a novel benchmark designed to evaluate Large Reasoning Models’ (LRMs) performance on simple tasks that favor intuitive system 1 thinking rather than deliberative system 2 reasoning. While LRMs have achieved significant breakthroughs in complex reasoning tasks through explicit chains of thought, their reliance on deep analytical thinking may limit their system 1 thinking capabilities. Moreover, a lack of benchmark currently exists to evaluate LRMs’ performance in tasks that require such capabilities. To fill this gap, S1-Bench presents a set of simple, diverse, and naturally clear questions across multiple domains and languages, specifically designed to assess LRMs’ performance in such tasks. Our comprehensive evaluation of 22 LRMs reveals significant lower efficiency tendencies, with outputs averaging 15.5 times longer than those of traditional small LLMs. Additionally, LRMs often identify correct answers early but continue unnecessary deliberation, with some models even producing numerous errors. These findings highlight the rigid reasoning patterns of current LRMs and underscore the substantial development needed to achieve balanced dual-system thinking capabilities that can adapt appropriately to task complexity [^1] The code and benchmark can be found in https://github.com/WYRipple/S1_Bench..
@@ 1 Introduction
Recent advances in Large Reasoning Models (LRMs), notably OpenAI’s o1/o3 OpenAI (2024) and the DeepSeek-R1 Guo et al. (2025) series, have propelled the development of Large Language Models (LLMs) by explicitly generating external chain-of-thought (COT) Wei et al. (2022) before producing final answers, achieving remarkable performance on complex reasoning tasks. Unlike traditional LLMs that primarily rely on intuitive, heuristic system 1 thinking, LRMs enhance their performance by employing analytical, deliberative system 2 thinking Qu et al. (2025); Li et al. (2025b), which includes sophisticated strategies such as self-reflection, multi-path exploration Li et al. (2025a); Yeo et al. (2025). Despite the benefits of system 2 thinking, a rigid reliance on this cautious thinking mode can also introduce new limitations, such as performance degradation in specific contexts Feng et al. (2025); Zhao et al. (2025) or overthinking Chen et al. (2024); Kumar et al. (2025a).
However, a research gap exists in thoroughly investigating how LRMs’ over-reliance on system 2 thinking affects their performance when confronting extremely simple questions better suited for intuition-driven system 1 processing. Specifically, existing research primarily analyzes LRM behaviors in complex tasks that naturally align with deliberative system 2 thinking, such as MATH for math problems Luo et al. (2025); Yang et al. (2025); Chen et al. (2024) or IFEval for instruction following Zhao et al. (2025). Although some benchmarks involve relatively simple questions for human Miao et al. (2021); Yan et al. (2025); Chiang and Lee (2024), these tasks are often adversarial or non-trivial for LLMs, requiring moderate reasoning and can not be effectively solved through intuitive system 1 processing. Thus, a benchmark to assess the system 1 thinking capability of LRMs is still lacking, further hindering our understanding of LRMs’ cognitive flexibility between the two systems Ziabari et al. (2025); Qu et al. (2025).
@@
To fill this research gap, we introduce the System 1 Thinking Capability Benchmark (S1-Bench), which measures the performance of LRMs across various simple tasks that commonly encountered in real-world applications. S1-Bench has the following three characteristics: (1) Simple. The questions are not hard for humans and can be easily answered by LLMs. LLMs with 7-9B parameters can robustly provide correct answers through direct responses when sampled across multiple temperatures. (2) Diverse. S1-Bench is not limited to simple reasoning problems; it encompasses four major categories and 28 subcategories in two languages (English and Chinese), including commonsense knowledge, loosely constrained instruction following, and straightforward analytical problems. (3) Natural. The questions are clear, without any misleading elements or ambiguities, ensuring they can be answered intuitively.
@@
S1-Bench conducted extensive evaluations across 22 LRMs, yielding the following key findings: (1) Current LRMs exhibit inefficiency and lack system 1 thinking capabilities, with average output lengths 15.5 times longer than small LLMs. (2) LRMs often arrive at correct answers early in their thinking process but continue unnecessary reasoning with gradually increasing content similarity. (3) Despite employing deep reasoning, LRMs exhibit accuracy degradation on simple questions compared to traditional LLMs. (4) LRMs show the ability to prejudge the simplicity of questions, yet they still exhibit inefficiency. These findings emphasize the significant distance LRMs must traverse to become powerful dual-system compatible models.
@@
  • To the best of our knowledge, S1-Bench is the first benchmark to evaluate the system 1 thinking capabilities of LRMs.
  • We introduce a workflow for constructing simple dataset suitable for system 1 evaluation.
  • Extensive experiments reveal the inefficiency and under-accuracy of LRMs on simple questions that are suited for system 1 thinking.
@@ 2.1 Large Reasoning Models
Large Reasoning Models (LRMs), characterized by explicitly generating external thinking processes before final answers Kumar et al. (2025b); Chen et al. (2025), achieve a paradigm shift from intuitive system 1 thinking to deliberative system 2 reasoning compared to traditional LLMs Li et al. (2025b); Qu et al. (2025), thus achieving superior performance on complex tasks. The development of recent LRMs has largely followed two main approaches: large-scale reinforcement learning (RL) and model distillation. Models trained via large-scale RL Guo et al. (2025); Team (2025b); Team et al. (2025) leverage reward-based optimization to gradually incentivize deliberative reasoning. In contrast, distillation-based LRMs OpenAI (2024); Min et al. (2024); Team (2025a); Ye et al. (2025); Muennighoff et al. (2025) acquire such abilities by transferring structured reasoning patterns from advanced teacher models.
@@ 2.2 Limitations of LRMs
While LRMs have shown significant performance gains through deliberate reasoning, rigid adherence to this overly cautious thinking can introduce new limitations. On one hand, intermediate reasoning steps can cause excessive token generation and unnecessary solving attempts Chen et al. (2024); Hashemi et al. (2025); Kumar et al. (2025a). On the other hand, LRMs’ performance can drop in specific contexts like safety scenarios Jiang et al. (2025); Zhao et al. (2025) and role-playing Feng et al. (2025). However, prior studies mainly evaluated LRMs on complex tasks that more suited for deliberative system 2 thinking. Our work examines how deliberative reasoning impacts extremely simple problems better matched to intuition-driven system 1 processing.
@@ 3.1 How to Ensure Simplicity?
We ensure questions are simple and suitable for system 1 thinking through two aspects.
@@ 3.1.1 A Priori Simplicity Constraints
We require all questions and answers to adhere to the following guidelines:
(1) Questions must be naturally, clearly expressed, and unambiguous, free of intentional traps.
(2) Answers must be unique or easily falsifiable (such as providing a three-letter English word).
We’ve further implemented category-specific constraints to ensure simplicity: Rsn: Limited to problems solvable with minimal reasoning or intuition. Kno: Restricted to common knowledge with unique, verifiable answers from sources like Wikipedia. If: Involve straightforward instructions without strict formatting requirements. Ana: Limited to questions whose answers can be directly inferred from the prompt, such as binary classification. These constraints ensure all questions remain straightforward for human respondents.
@@ 3.1.2 A Posteriori Simplicity Verification
Due to the biases existing between language models and humans Gallegos et al. (2024), questions that are simple for humans may be difficult for LLMs. Therefore, we introduce additional posteriori verification to ensure that questions are simple enough to be correctly and robustly answered by smaller LLMs from different families. This alleviates cognitive biases in traditional LLMs for system 1 thinking problems Li et al. (2025b).
@@ 3.2 Construction Workflow
To implement a priori constraints, we introduce data generators [^2] We select Claude-3.7-Sonnet and Qwen2.5-72B-Instruct. and quality discriminators [^3] We select GPT-4o and DeepSeek-V3-241226. that adhere to the a priori settings. To ensure posterior verification, we introduce simplicity validators [^4] We select four small LLMs: Qwen2.5-7B, Llama3.1-8B, Mistral-8B, and Gemma2-9B. The full model IDs are detailed in Table 4. and correctness evaluator [^5] We select GPT-4o.. The generators, discriminators, and evaluator are uniformly configured with temperature 0.7 and top-p=0.9, while validators are set with temperatures of 0.0, 0.2, and 0.4, each sampling 10 times. Additionally, we recruit three experienced graduate students as annotators, familiar with LLMs and fully comprehending the construction objectives of S1-Bench. For the prompts used in the workflow, refer to Table 11.
Subcategory Preparation. Initially, to ensure diversity, we draw on existing benchmark classification methods (e.g., MMLU, IFEval and GSM8K) and evaluation surveys Chang et al. (2024), selecting or merging existing subcategories, and designing new ones. We ensure that all subcategories meet the requirements for simplicity and provide an example question for each.
Generation and Modification. We use two data generators to create initial questions and answers for each candidate subcategory. The generation process adheres to prompt requirements (refer to Table 11), ensuring bilingual content that aligns with subcategory definitions while maintaining diverse perspectives. Each question is accompanied by auxiliary answers, with 50 initial bilingual question-answer pairs generated per subcategory. These pairs are independently assessed by three annotators and two quality discriminators , who ensure that the questions are unambiguous, have clear boundaries, and admit only one correct answer. For each question-answer pair, five independent assessments are conducted. Based on these evaluations, annotators discuss and collectively decide whether to retain, modify, or discard each question, ensuring alignment with the requirements outlined in Section 3.1.1. Retained questions are submitted to validators for testing, and an evaluator assesses the accuracy of responses. Questions with correct answers are included in S1-Bench, while others undergo reprocessing.
Questions with errors in any of the 30 sampling iterations undergo an iterative difficulty reduction process. These questions return to generators with specific prompts to decrease complexity, followed by the same discrimination, discussion, and evaluation procedures. Questions failing to achieve error-free performance after three difficulty reduction iterations are excluded from the workflow. The final S1-Bench comprises questions that satisfy both human-established a priori simplicity constraints and a posteriori simplicity verification.
@@
Llama3.3-70B 100.00 99.76 99.76 53.71
DeepSeek-v3 100.00 100.00 100.00 79.53
Table 1: Average accuracy and response token count of different LLMs, each sampled 10 times at three temperature settings.
@@ 3.3 Benchmark Statistics
S1-Bench comprises 422 question-answer pairs across four major categories and 28 subcategories, balanced with 220 English and 202 Chinese questions. We present the category distribution and descriptions with examples for all subcategories, which can be found in Figure 8 and Table 15,16 in the appendix. Figure 3 shows the token length distribution, with questions averaging 14.46 tokens.
To validate the absence of response bias from validators, we tested S1-Bench on five LLMs that serve as base models for commonly used LRMs. As shown in Table 1, these models achieved an accuracy exceeding 99% under three temperature settings, confirming the simplicity and robustness of S1-Bench.
Model ID Size Loose Format Strict Format L-Corr $\uparrow$ S-Corr $\uparrow$ $\mathtt{ART}\downarrow$
@@ 4.1 Baseline Models and Configurations
We evaluated 22 different LRMs, which were explicitly trained to first respond with a thinking process, and then generate a final answer. These LRMs include open-source families, such as DeepSeek Guo et al. (2025), Qwen Team (2025b), Nemotron Chris Alexiuk and Patel (2025), Light-R1 Wen et al. (2025), s1.1 Muennighoff et al. (2025), EXAONE Research et al. (2025), and Sky-T1 Griggs et al. (2025), as well as closed-source Hunyuan-T1 Tencent (2025), spanning from tiny (1.5B) to large (671B) parameter sizes [^6] More details in Appendix Table 5 (includes training data scale and algorithms).. Notably, OpenAI’s o-series models are not included as they do not disclose thinking processes to users. For each model, we consider two sets of generation configurations: Greedy sampling with temperature t = 0; Top-p sampling with temperature t = 0.6, top-p = 0.95 and sampling size k=5.
@@ 4.2 Evaluation Metrics
Generally, the output of LRMs should be separated by an “end thinking marker” (e.g., </think>) that distinguishes the thinking process from the final answer. However, LRMs do not always generate responses in correct format. To assess this, we calculate the percentage of responses that meet the required formatting criteria, categorized as follows (for top-p sampling setting, each metric is averaged across 5 generations):
S-Corr (Strict Format Correctness Rate): The percentage of responses that strictly adhere to the format, meaning they only contain one "end thinking marker" followed by a non-empty final answer.
L-Corr (Loose Format Correctness Rate): The percentage of responses that loosely adhere to the format, meaning they meet the format requirements except for responses with endless repetitions [^7] Detailed format analyses are in the Appendix C..
@@
Column Avg - 810.6 771.2 896.8 590.2 736.0 679.8 521.6 742.0 447.1 563.0 652.8
Improvement - $\times$10.9 $\times$18.5 $\times$109.7 $\times$11.0 $\times$14.9 $\times$14.1 $\times$13.8 $\times$62.7 $\times$14.8 $\times$16.7 $\times$15.5
Table 3: $\mathtt{ART}$ on the S1-bench across two languages and four main categories. Bold teal marks best performance, teal second best, bold burgundy worst, and burgundy second worst. Bold represents the maximum Improvement value for each language.
@@
We calculate the Average Response tokens (ART), which represent the average number of tokens in responses that meet the loose formatting requirements. Token counts are obtained using the Qwen2.5 tokenizer.

### v3

@@ Abstract
We introduce S1-Bench, a novel benchmark designed to evaluate the performance of Large Reasoning Models (LRMs) on simple tasks that favor intuitive system 1 thinking rather than deliberative system 2 reasoning. While LRMs have achieved significant breakthroughs in complex reasoning tasks through explicit chains of thought, their heavy reliance on system 2 thinking may limit their system 1 thinking capabilities. However, there is a lack of an appropriate benchmark for evaluating LRM’s system 1 thinking capabilities. To fill this gap, S1-Bench introduces a suite of simple, diverse, and natural questions across multiple domains and languages, specifically designed to assess LRMs’ performance on questions more suitable for system 1. We conduct extensive evaluations across 28 LRMs, revealing their inefficiency, inadequate accuracy, and limited robustness when handling simple questions. Additionally, we observe a gap between their difficulty perception and generation length. Overall, this work paves the way toward dual-system compatibility in the development of LRMs [^1] The code and benchmark can be found in https://github.com/WYRipple/S1_Bench..
@@ 1 Introduction
Recent advances in Large Reasoning Models (LRMs), notably OpenAI’s o1/o3 OpenAI (2024) and the DeepSeek-R1 Guo et al. (2025) series, have propelled the development of Large Language Models (LLMs). Unlike traditional LLMs that exhibit intuitive, heuristic system 1 thinking, LRMs demonstrate deliberate and analytical system 2 reasoning Qu et al. (2025a); Li et al. (2025b) by explicitly generating external chain-of-thought (COT) Wei et al. (2022) before producing final answers. Through sophisticated strategies such as self-reflection and multi-path exploration Li et al. (2025a); Yeo et al. (2025), LRMs can achieve strong performance in tasks that require system 2 thinking, including advanced mathematical and competition-level problems Yang et al. (2025a).
Benchmark Cross Realistic Multi- Acc.
Domain Scenarios lingual
AIME 6.67
GPQA 24.94
Olympiad-Bench 27.94
AMC 31.88
MATH 58.30
MMLU 66.27
GSM8K 87.45
ASDIV 97.51
GSM8K-zero 77.98
RoR-Bench 14.24
S1-Bench (ours) 100.00
Table 1: Characteristics of S1-Bench, with “Acc.” representing the average accuracy of four 7-9B LLMs. See Appendix A.1 for more details.
However, there remains a lack of appropriate benchmarks for evaluating LRMs’ system 1 thinking capabilities. Not all real-world problems require system 2 reasoning. The capacity to dynamically identify simple questions and address them efficiently contributes to both resource optimization and improved user satisfaction for LRMs. Nevertheless, current benchmarks either overemphasize difficulty, are simple yet lack domain diversity, or are only not hard for humans but involve unrealistic adversarial designs. Table 1 presents a collection of recent benchmarks aimed at mitigating the overthinking in LRMs Sui et al. (2025). The majority of these benchmarks are of high difficulty. For example, AIME and GPQA MAA Committees ; Rein et al. (2024) achieve less than 30% accuracy on conventional small LLMs, which are inherently more suitable for system 2 reasoning. Although some simple mathematical benchmarks are easy enough, such as GSM8K and ASDIV Cobbe et al. (2021); Miao et al. (2021), they often suffer from limited domain variety. Furthermore, some tasks that pose little challenge to humans but incorporate adversarial elements tend to lack relevance to realistic scenarios, such as GSM8K-zero Chiang and Lee (2024), which includes the correct answer in the questions. Thus, a benchmark to assess the system 1 thinking capability of LRMs is still lacking, further hindering our understanding of LRMs’ cognitive flexibility between the two systems Ziabari et al. (2025); Qu et al. (2025a).
@@
To fill this research gap, we introduce the System 1 Thinking Capability Benchmark (S1-Bench), which measures the performance of LRMs across various simple tasks that commonly encountered in real-world applications. S1-Bench has the following three characteristics: (1) Simple. The questions are not hard for humans and can be easily answered by LLMs. LLMs with 7-9B parameters can robustly provide correct answers through direct responses when sampled across multiple temperatures. (2) Diverse. S1-Bench is not limited to simple reasoning problems; it encompasses four major categories and 28 subcategories in two languages (English and Chinese), including reasoning problems, commonsense knowledge, instruction following, and analytical problems. (3) Natural. The questions are clear, without any misleading elements or ambiguities, ensuring they can be answered intuitively.
@@
We conduct extensive evaluations on S1-Bench across 28 LRMs, yielding the following key findings: (1) Current LRMs exhibit inefficiency and lack system 1 thinking capabilities across all types of questions, with average output lengths 15.5 times longer than small LLMs on S1-Bench. (2) Despite employing deep reasoning, several LRMs exhibit under-accuracy and limited robustness on simple questions. (3) LRMs exhibit “gut moment” at the beginning of some reasoning processes, showing gut feelings about task difficulty. Yet, even when recognizing a question’s simplicity, LRMs often fail to produce shorter responses—revealing a gap between their difficulty awareness and generation behavior. These findings emphasize the significant distance LRMs must traverse to become powerful dual-system compatible models.
@@
  • To the best of our knowledge, S1-Bench is the first benchmark to evaluate the system 1 thinking capabilities of LRMs, which paves the way for dual-system compatibility.
  • We introduce a workflow for constructing a simple dataset for system 1 evaluation.
  • Extensive experiments reveal the inefficiency, under-accuracy, and limited robustness of LRMs on simple questions.
  • We find that LRMs exhibit “gut moment” on simple problems, and reveal a gap between their difficulty perception and generation length.
@@ 2.1 Large Reasoning Models
Large Reasoning Models (LRMs), characterized by explicitly generating external thinking processes before final answers Kumar et al. (2025b); Chen et al. (2025a), achieve a paradigm shift from intuitive system 1 thinking to deliberative system 2 reasoning compared to traditional LLMs Li et al. (2025b); Qu et al. (2025a), thus achieving superior performance on complex tasks. The development of recent LRMs has largely followed two main approaches: large-scale reinforcement learning (RL) and model distillation. Models trained via large-scale RL Guo et al. (2025); Team (2025b); Team et al. (2025b) leverage reward-based optimization to gradually incentivize deliberative reasoning. In contrast, distillation-based LRMs OpenAI (2024); Min et al. (2024); Team (2025a); Ye et al. (2025); Muennighoff et al. (2025); Zhang et al. (2025b) acquire such abilities by transferring structured reasoning patterns from advanced teacher models.
@@ 2.2 Limitations of LRMs
While LRMs have shown significant performance gains through deliberate reasoning, rigid adherence to this overly cautious thinking can introduce new limitations. On one hand, intermediate reasoning steps can cause excessive token generation and unnecessary solving attempts Chen et al. (2024b); Hashemi et al. (2025); Kumar et al. (2025a), even leading to redundancy in the hidden layers Chen et al. (2024c, 2025b). On the other hand, LRMs’ performance can drop in specific contexts like safety scenarios Jiang et al. (2025) and role-playing Feng et al. (2025). However, prior studies mainly evaluated LRMs on complex tasks that more suited for deliberative system 2 thinking. Our work examines how deliberative reasoning impacts extremely simple problems better matched to intuition-driven system 1 processing.
@@ 3.1 How to Ensure Simplicity?
We ensure questions are simple and suitable for system 1 thinking through the following two aspects.
@@ 3.1.1 A Priori Simplicity Constraints
We begin by generating question–answer pairs through collaboration between humans and LLMs. Each pair is required to satisfy both the general and the category-specific simplicity criteria.
The general simplicity criteria requires that: (1) Questions must be naturally and clearly expressed, unambiguous, and free of intentional traps. (2) Answers must be unique or easily falsifiable (e.g., providing a three-letter English word).
The category-specific simplicity criteria are as follows. Rsn: Limited to problems solvable with minimal reasoning or intuition. Kno: Restricted to common knowledge with unique, verifiable answers from sources like Wikipedia. If: Involve straightforward instructions without strict formatting requirements. Ana: Limited to questions whose answers can be directly inferred from the prompt, such as binary classification. These constraints ensure all questions remain straightforward for human respondents.
@@ 3.1.2 A Posteriori Simplicity Verification
Due to the biases existing between language models and humans Gallegos et al. (2024), questions that are simple for humans may be difficult for LLMs. Therefore, we introduce additional posteriori verification to ensure that questions are simple enough to be correctly and robustly answered by smaller LLMs from different families.
@@ 3.2 Construction Workflow
Subcategory Preparation. To ensure diversity, we refer to the subcategories included in existing benchmarks (e.g., MMLU, IFEval, and GSM8K) and evaluation surveys Chang et al. (2024) to select, merge, or design subcategories for S1-bench, ensuring that each meets the simplicity requirements. The definition and example question for each subcategory can be found in Appendix A.2.
Implementation of A Priori Simplicity. First, we use two data generators [^2] We select Claude-3.7-Sonnet and Qwen2.5-72B-Instruct. to create 100 initial bilingual question-answer pairs for each candidate subcategory. The data generation prompt (see Appdiix A.3) explicitly incorporates the subcategory definitions, along with both the general and category-specific simplicity criteria, while also aiming to ensure diversity in the generated questions. Second, these question–answer pairs are then independently evaluated by three annotators and two quality discriminators [^3] We select GPT-4o and DeepSeek-V3-241226. according to the general and category-specific simplicity criteria (see Appdiix A.3 for prompt of discriminators), resulting in five evaluation outcomes per pair. The three annotators are experienced graduate students familiar with LLMs and well-acquainted with the goals of S1-Bench. Finally, based on these evaluation outcomes, three annotators discuss and collectively decide whether to retain, modify, or discard each question.
Implementation of A Posteriori Simplicity. First, each question obtained from the previous stage is input into the small LLM validators [^4] We select four small LLMs: Qwen2.5-7B, Llama3.1-8B, Mistral-8B, and Gemma2-9B. The full model IDs are detailed in Table B.1. with 7~9 B parameters. For each question, we sample 10 answers at three different temperature settings (0, 0.2, and 0.4), resulting in a total of 30 responses per question. These responses are then individually evaluated for correctness using GPT-4o. Second, if all 30 sampled responses are correct, the question is accepted into S1-Bench. Otherwise, the question is returned to the generators, where a difficulty-reduction prompt (see Appendix 10) is applied to simplify it. The simplified questions then undergoes the same subsequent process. Finally, questions fail to meet the full-accuracy criterion (i.e., 30 out of 30 correct) after three rounds of difficulty reduction are excluded from the workflow.
The final S1-Bench comprises questions that satisfy both human-based a priori simplicity constraints and LLM-based a posteriori simplicity verification.
@@
Qwen3-32B (w/o think) 100.00 100.00 100.00 103.30
Qwen3-14B (w/o think) 100.00 100.00 100.00 86.35
Qwen3-8B (w/o think) 100.00 100.00 99.76 90.54
Qwen3-1.7B (w/o think) 98.10 97.16 95.73 114.32
Table 2: Average accuracy (acc@k) and response token count of different LLMs, each sampled 10 times at three temperature settings.
@@ 3.3 Benchmark Statistics
S1-Bench comprises 422 question-answer pairs across four major categories and 28 subcategories, balanced with 220 English and 202 Chinese questions. Figure 2 shows the token length distribution, with questions averaging 14.46 tokens. To ensure that the a posteriori verification process does not introduce simplicity only tailored to the small LLM validator, we evaluate S1-Bench on five additional LLMs and on Qwen3 Family with reasoning modes disabled. As shown, even the 1.7B model achieves over 98% accuracy at temperature 0.
Model ID Size Loose Format Strict Format L-Corr $\uparrow$ S-Corr $\uparrow$ Tokens $\downarrow$
@@ 4.1 Baseline Models and Configurations
We evaluated 28 different LRMs, which are explicitly trained to first respond with a thinking process, and then generate a final answer. These LRMs include open-source families, such as DeepSeek Guo et al. (2025), Qwen Yang et al. (2025a), Nemotron Bercovich et al. (2025), Light-R1 Wen et al. (2025), s1.1 Muennighoff et al. (2025), EXAONE Research et al. (2025), and Sky-T1 Griggs et al. (2025), as well as closed-source Hunyuan-T1 Tencent (2025), spanning from tiny (1.5B) to large (671B) parameter sizes [^5] Model details are presented in Appendix B.1.. Notably, OpenAI’s o-series models are not included as they do not disclose thinking processes to users. For each model, we consider two sets of generation configurations: Greedy sampling with temperature t=0; Top-p sampling with temperature t=0.6, top-p=0.95 and sampling size k=5. Only top-p sampling results are reported in the main text; greedy decoding results are provided in the Appendix C.1.
@@ 4.2 Evaluation Metrics
To assess the formatting quality of LRM responses, we compute the proportion of responses that satisfy the following two formatting criteria (averaged over 5 runs for top-p sampling).
S-Corr (Strict Format Correctness Rate): In general, an end thinking marker (e.g., </think>) is expected to separate the thinking process from the non-empty final answer. S-Corr measures the proportion of responses that satisfy this criterion.
L-Corr (Loose Format Correctness Rate): LRMs may occasionally generate responses with endless thinking. L-Corr quantifies the proportion of responses that do not exhibit this failure mode.
Detailed format types are given in Appendix B.4.
@@
Column Avg - 809.1 766.0 785.1 591.5 718.4 695.8 545.9 677.9 482.0 576.7 650.3
Improvement - $\times$10.8 $\times$18.4 $\times$96.0 $\times$11.1 $\times$14.5 $\times$14.4 $\times$14.5 $\times$57.3 $\times$16.0 $\times$17.1 $\times$15.5
Table 4: Average response tokens in the top-p sampling setting on the S1-bench across two languages and four main categories. Bold teal marks best performance, teal second best, bold burgundy worst, and burgundy second worst. Bold represents the maximum Improvement value for each language.
@@
We calculate the average token counts for responses (Tokens) except for those generate endless thinking. Token counts are obtained using the Qwen2.5 tokenizer.